# Spec2VolCAMU-Net: A Spectrogram-to-Volume Model for EEG-to-fMRI Reconstruction based on Multi-directional Time-Frequency Convolutional Attention Encoder and Vision-Mamba U-Net

### v1

@@ 4 Results and Discussion
Figure 4: Visual comparison of reconstructed fMRI volumes from Spec2VolCAMU-Net and certain baseline methods on the NODDI dataset. The first column shows the ground truth (GT) fMRI volume, while the subsequent columns display the results from E2fNet, E2fGAN, and Spec2VolCAMU-Net.
@@
Figure 5: Visual comparison of reconstructed fMRI volumes from Spec2VolCAMU-Net and certain baseline methods on the Oddball dataset. The first column shows the ground truth (GT) fMRI volume, while the subsequent columns display the results from E2fNet, E2fGAN, and Spec2VolCAMU-Net.
@@ 5 Conclusion
This research presents Spec2VolCAMU-Net, a novel EEG-to-fMRI reconstruction model employing a proposed Multi-directional Time-Frequency Convolutional Attention Encoder (MD-TF-CAE) and a Vision-Mamba U-Net decoder. The model effectively integrates local and global EEG features and efficiently models spatial dependencies, addressing limitations of prior methods. Evaluated on NODDI, Oddball, and CN-EPFL datasets, Spec2VolCAMU-Net achieves state-of-the-art SSIM scores, significantly improving structural fidelity in reconstructed fMRI volumes, and demonstrates competitive PSNR, balancing structural preservation and detail accuracy. While showing robust performance, limitations include potential for enhanced high-frequency detail restoration and sensitivity to severe EEG artifacts, suggesting future work on improved noise modeling and artifact rejection. Extending validation to more complex tasks and populations, and exploring advanced techniques like multimodal consistency regularization, will be crucial for broadening the model’s applicability as a cost-effective neuroimaging tool.
@@ Acknowledgments
This work is supported by the Scientific and Technological Research Program of the Chongqing Education Commission (KJQN202201142), the Natural Science Foundation of Chongqing (2024NSCQ-MSX1731), the Chongqing University of Technology Research and Innovation Team Cultivation Program (2023TDZ012), the Chongqing Municipal Key Project for Technology Innovation and Application Development (CSTB2024TIAD-KPX0042), and the National Natural Science Foundation of P.R. China (61173184).

### v2

@@ 4.2 Qualitative Visual Comparisons
Figure 5: Visual comparison of reconstructed fMRI volumes from Spec2VolCAMU-Net and certain baseline methods on the NODDI dataset. The first column shows the ground truth (GT) fMRI volume, while the subsequent columns display the results from E2fNet, E2fGAN, and Spec2VolCAMU-Net. The brain regions that improve the most with Spec2VolCAMU-Net are highlighted with red boxes.
@@
Figure 6: Visual comparison of reconstructed fMRI volumes from Spec2VolCAMU-Net and certain baseline methods on the Oddball dataset. The first column shows the ground truth (GT) fMRI volume, while the subsequent columns display the results from E2fNet, E2fGAN, and Spec2VolCAMU-Net. The brain regions that improve the most with Spec2VolCAMU-Net are highlighted with red boxes.
@@
4.3 Results of Ablation Study
To dissect the contribution of each key component within the Spec2VolCAMU-Net architecture, we conducted a comprehensive ablation study on the CN-EPFL dataset. The results, detailed in Table 3, systematically quantify the impact of the VM-UNet decoder, the MHSA module, and the individual convolutional paths within the MD-TF-CAE.
Impact of the Vision-Mamba U-Net Decoder: The most significant performance degradation was observed when the VM-UNet decoder was replaced with a standard U-Net architecture, composed of conventional convolutional and up-sampling layers. This change resulted in a substantial drop in SSIM from 0.788 to 0.741 and a decrease in PSNR from 23.837 to 23.112. This finding strongly validates the critical role of the VM-UNet in reconstructing high-fidelity fMRI volumes. A standard U-Net, while effective at capturing hierarchical features through its skip connections, is constrained by the local receptive field of its convolutional kernels. It struggles to explicitly model the long-range spatial dependencies that characterize distributed brain networks. In contrast, the VSS blocks within the VM-UNet, with their underlying state-space mechanism and 2D selective scan, can efficiently capture global contextual information and model complex, non-local interactions across the entire 3D volume. This capability is paramount for generating anatomically coherent and structurally accurate brain maps, as reflected by the large performance gap.
Impact of the Multi-Head Self-Attention in the Encoder: Removing the MHSA module from the MD-TF-CAE led to the second-largest performance drop, with SSIM falling to 0.755 and PSNR to 23.345. This underscores the importance of capturing global relationships within the EEG spectrogram before decoding. While the multi-directional convolutions excel at extracting local time-frequency patterns, the MHSA module integrates this information by attending to relationships across all channels, time steps, and frequency bins. This allows the encoder to form a holistic representation that captures how activity in one electrode/frequency band relates to another, which is neurophysiologically crucial for inferring distributed fMRI activation. Without MHSA, the model relies solely on the convolutional receptive fields to aggregate information, leading to a less comprehensive encoding and subsequently poorer reconstruction quality.
Impact of Multi-directional Convolutional Paths: Ablating the individual convolutional paths within the encoder—temporal, frequency, and joint time-frequency—resulted in more modest but still significant performance decreases. Removing the frequency convolution path caused the most notable drop among the three (SSIM to 0.765), suggesting that accurately modeling spectral patterns is particularly vital for the EEG-fMRI mapping. The degradation from removing the temporal (SSIM to 0.769) and joint paths (SSIM to 0.772) confirms their utility in capturing dynamic signal changes and their interactions, respectively. These results collectively demonstrate the benefit of the multi-path design. By processing the spectrogram from different perspectives in parallel, the encoder builds a more robust and comprehensive feature set than a single, monolithic convolutional block could. The fact that the model’s performance degrades with the removal of any single path confirms that they learn complementary, rather than entirely redundant, features.
In conclusion, the ablation study provides compelling empirical evidence for our architectural choices. The results highlight a clear hierarchy of importance: the long-range spatial modeling of the VM-UNet decoder is most critical, followed by the global feature integration of the MHSA in the encoder, and complemented by the rich, multi-faceted local feature extraction of the directional convolutional paths. Each component makes a synergistic contribution, and their combined effect is what enables Spec2VolCAMU-Net to achieve its state-of-the-art performance.
Table 3: Ablation study results on the CN-EPFL dataset.
MD-TF-CAE Encoder VM-UNet Decoder SSIM PSNR
Temporal Conv. Frequency Conv. Joint Time-Freq. Conv. MHSA
0.788 23.837
$\times$ 0.766 23.185
$\times$ 0.778 23.665
$\times$ 0.787 23.801
$\times$ $\times$ 0.787 23.825
$\times$ 0.786 23.481
$\times$ $\times$ 0.786 23.490
$\times$ 0.769 23.124
$\times$ $\times$ 0.775 23.461
The best results are highlighted in red.
4.4 Reasons for the Superior Performance
@@ 4.5 Limitations and Future Work
A key limitation is that the current model and its evaluation focus on the reconstruction of individual fMRI volumes at discrete time points. Consequently, the employed metrics, SSIM and PSNR, primarily assess spatial fidelity rather than the temporal coherence of the BOLD signal. To enhance the model’s relevance and persuasiveness for the neuroimaging community, particularly for applications like functional connectivity analysis, a critical future direction is to validate the temporal dynamics of the reconstructed fMRI signal. We plan to address this by extending the framework to generate entire BOLD time series. The quality of these series will be evaluated by comparing them against ground truth data using metrics such as the Pearson correlation coefficient of time courses extracted from predefined Regions of Interest (ROIs) or by computing pseudo R-squared values. Demonstrating that the reconstructed voxel time series maintain high temporal consistency and faithfully capture the underlying neural dynamics is a crucial next step that we will emphasize in our future work.
@@ 5 Conclusion
In this paper, we presented Spec2VolCAMU-Net, a novel deep learning architecture that significantly advances the state-of-the-art in synthesizing high-fidelity 3D fMRI volumes from EEG spectrograms. The model’s strength lies in its synergistic design, featuring a Multi-directional Time-Frequency Convolutional Attention Encoder (MD-TF-CAE) to capture a rich hierarchy of neuroelectric features and a Vision-Mamba U-Net decoder to efficiently model the long-range spatial dependencies inherent in brain activation maps. Our comprehensive evaluations across three diverse public datasets confirmed its superiority, establishing new benchmarks in structural similarity (SSIM) while maintaining competitive pixel-level accuracy (PSNR).
Crucially, we position the success in high-quality single-volume reconstruction as a foundational and indispensable first step. The ultimate ambition of EEG-to-fMRI synthesis is to generate temporally coherent fMRI time series suitable for dynamic brain analysis. We argue that ensuring the high spatial fidelity of each individual volumetric snapshot is a prerequisite for faithfully modeling the temporal evolution of neural activity. Therefore, this work establishes the robust spatial groundwork necessary to tackle the subsequent, more complex challenge of achieving full and accurate spatio-temporal reconstruction.
Building upon this validated spatial foundation, our future research will prioritize extending the model to generate and validate complete BOLD time series, with a focus on their temporal consistency. By continuing to refine this technology, Spec2VolCAMU-Net and its successors hold the promise of bridging the gap between high-temporal-resolution EEG and high-spatial-resolution fMRI, ultimately paving the way for more accessible, cost-effective, and powerful functional neuroimaging tools for both clinical and research communities.
@@ Acknowledgments
This research was funded by the Scientific and Technological Research Program of the Chongqing Education Commission (KJZD-K202303103, KJZD-K202501107, KJQN202501104), the Natural Science Foundation of Chongqing (CSTB2024NSCQ-MSX0118), the Chongqing Municipal Key Project for Technology Innovation and Application Development (CSTB2024TIAD-KPX0042, CSTB2025TIAD-KPX0002), the College Student Innovation and Entrepreneurship Project (2025CXXL025), and the National Natural Science Foundation of P.R. China (61173184).